- Back C - K ZH-C – H C=C-Rname of this reaction?
OH +®, CH-CEC-R
Answer #1

-e + :C=c-R tren СЕС-К lue door § 6-H C=C-R The above reaction is a mucleoprile addition at reaction in which Alkyne will act